# GA-SmaAt-GNet: Generative Adversarial Small Attention GNet for Extreme Precipitation Nowcasting

### v1

@@ 5.3 Activation heatmaps
Explaining the outcomes of deep learning models is essential, and following the approach in sar , we use Grad-CAM grad to generate activation heatmaps. These heatmaps highlight the input regions influencing high values in the output, offering valuable insights into the decision-making process of the model. Originally designed for visual explanations in classification networks, Grad-CAM calculates the gradient of the class score with respect to feature map activations to determine the importance weights of neurons. It then combines these weights linearly with the forward activation maps, followed by ReLU, to produce the final outcome. To generate activation heatmaps for our nowcasting task, we convert it into an image segmentation task by binarizing the output using a threshold of 0.5 mm/h in the same way we did for our binary evaluation metrics.
@@
Examining the encoder activation heatmaps in Fig. 10(b), we notice distinct patterns. The precipitation map encoder’s heatmaps exhibit higher activation at depths 0 and 1 in regions corresponding to elevated precipitation levels in the ground truth. In contrast, the precipitation mask encoder’s heatmaps at depth 0 show increased activation in areas with higher precipitation in the model input, while at depth 1, it shows increased activation around the high precipitation area in the input. As we move to depths 2, 3, and 4, the activation areas for both encoders become more abstract and less directly linked to regions of higher precipitation, an observation also made by sar at lower levels of their architecture. When comparing both encoders at all depths, we see that they are generally activated in different areas, suggesting complementary contributions to the prediction. We can also observe the effects of the CBAM layers if we compare the double DSC activation maps and the CBAM activation maps. We can observe in which areas the CBAM layers emphasize and de-emphasize the double DSC activations. As an example, consider the precipitation map encoder at depth 1; after the application of the CBAM layer, the region corresponding to higher precipitation in the model’s prediction exhibits increased activation. Additionally, the influence of CBAM layers is more pronounced in the precipitation mask encoder than in the precipitation map encoder, evident from the greater differences between the activation heatmaps of the DSC blocks and the CBAM layers in the precipitation mask encoder.
@@ 6 Conclusion and future work
In this paper we introduced GA-SmaAt-GNet, a novel generative adversarial model for extreme precipitation nowcasting. GA-SmaAt-GNet uses two methodologies to more accurately nowcast extreme precipitation: the integration of precipitation masks in a novel SmaAt-GNet architecture and a generative adversarial approach with an attention-augmented discriminator, which resulted in a substantial advancement for extreme precipitation nowcasting. Furthermore, we demonstrated that both aleatoric and epistemic uncertainty increase with higher precipitation intensities, emphasizing the inherent challenges associated with the task of extreme precipitation nowcasting. Finally, we utilized Grad-CAM to generate activation heatmaps for various parts of our proposed GA-SmaAt-GNet, providing valuable insights into the model’s predictions.

### v2

@@ 5.3 Activation heatmaps
Explaining the outcomes of deep learning models is essential, and following the approach in [29], we use Grad-CAM [45] to generate activation heatmaps. These heatmaps highlight the input regions influencing high values in the output, offering valuable insights into the decision-making process of the model. Originally designed for visual explanations in classification networks, Grad-CAM calculates the gradient of the class score with respect to feature map activations to determine the importance weights of neurons. It then combines these weights linearly with the forward activation maps, followed by ReLU, to produce the final outcome. To generate activation heatmaps for our nowcasting task, we convert it into an image segmentation task by binarizing the output using a threshold of 0.5 mm/h in the same way we did for our binary evaluation metrics.
@@
Examining the encoder activation heatmaps in Fig. 11(b), we notice distinct patterns. The precipitation map encoder’s heatmaps exhibit higher activation at depths 0 and 1 in regions corresponding to elevated precipitation levels in the ground truth. In contrast, the precipitation mask encoder’s heatmaps at depth 0 show increased activation in areas with higher precipitation in the model input, while at depth 1, it shows increased activation around the high precipitation area in the input. As we move to depths 2, 3, and 4, the activation areas for both encoders become more abstract and less directly linked to regions of higher precipitation, an observation also made by [29] at lower levels of their architecture. When comparing both encoders at all depths, we see that they are generally activated in different areas, suggesting complementary contributions to the prediction. We can also observe the effects of the CBAM layers if we compare the double DSC activation maps and the CBAM activation maps. We can observe in which areas the CBAM layers emphasize and de-emphasize the double DSC activations. As an example, consider the precipitation map encoder at depth 1; after the application of the CBAM layer, the region corresponding to higher precipitation in the model’s prediction exhibits increased activation. Additionally, the influence of CBAM layers is more pronounced in the precipitation mask encoder than in the precipitation map encoder, evident from the greater differences between the activation heatmaps of the DSC blocks and the CBAM layers in the precipitation mask encoder.
@@ 6 Conclusion and future work
In this paper we introduced GA-SmaAt-GNet, a novel generative adversarial model for extreme precipitation nowcasting. GA-SmaAt-GNet uses two methodologies to more accurately nowcast extreme precipitation: the integration of precipitation masks in a novel SmaAt-GNet architecture and a generative adversarial approach with an attention-augmented discriminator. The validity and applicability of the proposed models are shown on real-life precipitation dataset of Netherlands spanning 25 years. Furthermore, we demonstrated that both aleatoric and epistemic uncertainty increase with higher precipitation intensities, emphasizing the inherent challenges associated with the task of extreme precipitation nowcasting. Finally, we utilized Grad-CAM to generate activation heatmaps for various parts of our proposed GA-SmaAt-GNet, providing valuable insights into the model’s predictions. While the methodologies presented in this paper showcase promising results for extreme precipitation nowcasting, further research is essential to refine the accuracy and certainty of predictions, particularly for longer lead times. For the future work one may consider adapting other possible loss functions such as weighted loss or incorporating other weather variables and exploring the inclusion of different data sources.